# A non-parametric approach for estimating consumer valuation distributions using second price auctions

### v2

@@ Abstract
We focus on online second price auctions, where bids are made sequentially, and the winning bidder pays the maximum of the second-highest bid and a seller specified reserve price. For many such auctions, the seller does not see all the bids or the total number of bidders accessing the auction, and only observes the current selling prices throughout the course of the auction. We develop a novel non-parametric approach to estimate the underlying consumer valuation distribution based on this data. Previous non-parametric approaches in the literature only use the final selling price and assume knowledge of the total number of bidders. The resulting estimate, in particular, can be used by the seller to compute the optimal profit-maximizing price for the product. Our approach is free of tuning parameters, and we demonstrate its computational and statistical efficiency in a variety of simulation settings, and also on an Xbox 7-day auction dataset on eBay.
@@ 1 Introduction
In a second price auction with reserve price, the product on sale is awarded to the highest bidder if the corresponding bid is higher than a seller-specified reserve price $r$. The price paid by the winner is however, the maximum of the reserve price and the second highest bid. These auctions have been the industry standard for a long time, and are attractive to sellers as they induce the bidders to bid their true “private value” for the product, i.e., the maximum price they wish to pay for it. While some platforms have recently moved to first-price auctions, the second-price auction is still widely used on E-commerce platforms such as eBay, Rokt and online ad exchanges such as Xandr. The analysis of data obtained from these auctions presents unique challenges. For a clear understanding of these challenges, we first discuss in detail the auction framework, the observed data and the quantity of interest that we want to estimate/extract.
@@
Auction framework: We consider an auction setting where a single product is on sale for a fixed time window $0,\tau]$. The seller sets the reserve price $r$, which is used as the current selling price at time 0. Any bidder who arrives subsequently is allowed to place a bid only if his/her bid value is higher than current selling price at that time. If the bid is placed, the current selling price is updated to the second-highest bid value among the set of all placed bids up to that time, including this latest bid (the reserve price is also treated as a placed bid).  [^1] Typically, a small increment (e.g. $\mathdollar 0.01$) is also added to the second highest bid, but this insignificant increment is unlikely to influence bidder’s behaviour, and we ignore it in our analysis for ease of exposition. For example, suppose the current selling price at a given time is $\mathdollar 4$ and the highest placed bid value up to that time is $\mathdollar 5$. If a bidder comes and bids $\mathdollar 3$, the bid will not be placed. If the bidder were to bid $\mathdollar 4.5$, the bid would be placed and the current selling price would be updated to $\mathdollar 4.5$ (since now the second largest placed bid is $\mathdollar 4.5$). If the bidder were to bid $\mathdollar 6$, the bid would be placed and the current selling price would be updated to $\mathdollar 5$. At the end of the auction period, if no bid above the reserve price is placed, the item goes unsold, otherwise it is sold to the highest bidder at the selling price at time $\tau$. This final selling price is the second highest placed bid (including the reserve price) throughout the course of the auction.
@@
Figure 1: An illustration of a single second price auction. True bid values are generated from a Pareto distribution and reserve price is at $\mathdollar 2$. Blue vertical lines are the bid values, black horizontal lines are the current selling prices, and black dots are the time points when the bids are made.
@@
An illustration of a single second price auction: Figure 1 provides a concrete illustration of how a second price auction works. The data for this auction has been simulated from a setting where the bids follow a Pareto distribution with location parameter $3$ and dispersion parameter $100$. The waiting times between bids are generated from an exponential distribution with rate parameter $1$. In total, $100$ bids are generated in a time period $\tau$ of around $115$ minutes. The reserve price for the auction is $\mathdollar 2$. In Figure 1, the bid values and the current selling prices during the course of the auction are represented by the blue vertical lines and the black horizontal lines, respectively. The black dots on the black horizontal lines represents the time points (in minutes) of the $100$ bids. As can be seen from Figure 1, the initial selling price is equal to the reserve price ($\mathdollar 2$). We have the first bid of around $\mathdollar 8.05$ at $0.55$ minutes. Since it’s higher than the reserve price $\mathdollar 2$, the reserve price still remains the current selling price at $0.55$ minutes. However, when the second bid of $\mathdollar 5.09$ occurs at $5.96$ minutes ($5.41$ minutes after the first bid’s occurrence), the current selling price jumps to $\mathdollar 5.09$ as it’s the second highest value among the reserve price ($\mathdollar 2$) and the two existing placed bids ($\mathdollar 8.05$ and $\mathdollar 5.09$). We don’t observe any jumps in the current selling price for the next few bids as they all are less than the current selling price of $\mathdollar 5.09$ (and hence are not placed). Then we see another jump at $9.65$ minutes, where a bid of $\mathdollar 12.82$ which makes the current selling price jump to $\mathdollar 8.05$. The next few bids again happen to be less than the current selling price ($\mathdollar 8.05$). At around $24$ minutes, we see a last jump in the current selling price to $\mathdollar 10.14$ (based on a new bid of $\mathdollar 10.14$ which exceeds $\mathdollar 8.05$). The subsequent bids are all less than $\mathdollar 10.14$, it remains the current selling price throughout the rest of the auction period. This can be observed through the flat horizontal black line at $\mathdollar 10.14$ in the time period of $(24.7,114.43)$ minutes. The final selling price for the auction is therefore $\mathdollar 10.14$. The observed data for the above auction is the sequence of current selling prices given by $(\mathdollar 2,\mathdollar 5.09,\mathdollar 8.05,\mathdollar 10.14)$ and the sequence of times at which there was a change in the current selling price, given by $(5.96,9.65,24)$.
@@
Observed data: The observed data is the sequence of current selling price values (also sometimes referred to as the standing price) throughout the course of the auction, and the times at which there is a change in the selling/standing price. Typically, such data is available for multiple auctions of the same product. For example, in Section 5, we analyze data with current selling prices for $93$ different eBay $7$-day auctions for Xbox. A key observation to make here is that consumers who access the auction but have bids which are less than the current selling price (standing price) are not allowed to place their bids. In other words, instead of observing the bids of all the customers who access the auction, we only observe the running second maximum of such bids.
@@
Figure 2: An illustration of 4 separate second price auctions for a given product. For each auction, true bid values are generated from a discrete Uniform distribution and reserve price is respectively set to $\mathdollar 10,\mathdollar 5,\mathdollar 13,\mathdollar 17$. Blue vertical lines are the bid values, black horizontal lines are the current selling prices, and black dots are the time points when the bids are made.
@@
An illustration of multiple second price auctions: Figure 2 provides a concrete illustration of a setting with multiple (four) second price auctions for the same product. The data for these auctions has been simulated from a setting where the bids follow a discrete uniform distribution with lower bound 1 and upper bound 25. The waiting times between bids are generated from an exponential distribution with rate parameter $1$, and the time period for each auction is 10 days. The reserve prices for the four auctions, respectively, are $\mathdollar 10,\mathdollar 5,\mathdollar 13,\mathdollar 17$. Just like in Figure 1, the bid values and the current selling prices during the course of each auction are represented by the blue vertical lines and the black horizontal lines, respectively. The black dots on the black horizontal lines represents the time points (in days) of the bids. We can see that for auctions in the upper left and upper right, the corresponding item is sold above the reserve price (at final selling price of $\mathdollar 19$ and $\mathdollar 25$ respectively). On the other hand, for the auction in the lower left, the item is sold at the reserve price, since there is only one bid placed above/at the reserve price. Finally, for the auction in the lower right, the item is not sold since no bid is placed above or at the reserve price. We will keep coming back to this integrated example to clarify and illustrate various concepts and notions that are introduced in the paper.
@@
Quantity of interest: Each bidder in the consumer population is assumed to have an independent private valuation (IPV) of the product. The IPV assumption in particular makes sense for products that are used for personal use/consumption (such as watches, jewelry, gaming equipment etc.) and is commonly used in the modeling of internet auctions (see Song (2004); Hou and Rego (2007); George and Hui (2012) and the references therein). Economic theory suggests that the dominant strategy for a bidder in a second-price auction is to bid one’s true valuation (Vickrey (1961)). The quantity that we want to estimate from the above data is the distribution of the valuations of the product under consideration for the consumer population. We refer to this as the consumer valuation distribution, and denote it by $F$. As noted in George and Hui (2012), knowledge of $F$ provides the consumer demand curve for the product, and hence can be used by the seller to identify the profit-maximizing price (see the discussion in Section 4.1 from George and Hui (2012)).
@@
This paper $2^{nd}$ Non-parametric Yes No No No
@@
Table 1 provides a comparison of the main relevant features of key related papers and the proposed methodology in the paper. As the table demonstrates (and to the best of our knowledge), none of the existing non-parametric methods for second price auctions use the entire sequence of current selling price values to estimate the consumer valuation distribution $F$. This was the primary motivation for our work, as only using the final selling price (and possibly the maximum placed bid, if available) leaves out a lot of available information. Including the current selling price information throughout the course of the auction however, involves significant conceptual and methodological challenges. As observed previously, the final selling price (second largest placed bid) is also the second largest (potential or placed) bid of all the consumers who accessed the auction. Hence, under relevant assumptions (see beginning of Section 2), the final selling price can be interpreted as the second largest order statistic of i.i.d. samples from $F$. This interpretation is central to the methodology developed in George and Hui (2012) and Song (2004). However, as the authors in George and Hui (2012) point out, the second largest current selling price throughout the course of the auction is not necessarily the third largest order statistic among all (placed or potential) bids, unless some severely restrictive and unrealisitc assumptions are imposed on the order in which bidders arrive in the auction. Hence, extending the methodology in George and Hui (2012) to include the entire sequence of current selling prices is not feasible; a completely new method of attack is needed.
@@
In this paper, we fill this gap in the literature and develop novel methodology for non-parametric estimation of the consumer valuation distribution in second-price ascending bid auctions which uses the entire sequence of current selling prices. Additionally, the total number of consumers accessing the auction is not assumed to be known (unlike George and Hui (2012)), and the highest placed bid in the auction is also not assumed to be known (unlike Backus and Lewis (2025)). Incorporating the above novel features is very challenging, and the methodological development (provided in Section 2) is quite involved. The extensive simulation results in Section 4 demonstrate the significant accuracy gains in the estimation of $F$ that can be obtained by using the entire sequence of current selling prices as opposed to just the final selling price. We note that incorporating the current selling prices during the entire auction does come with a cost. In particular, two additional assumptions (compared to those in George and Hui (2012)) that need to be made about the rate of bidder arrival and number of bids made by a single consumer. We find some deviations from these two assumptions in the eBay $7$-day Xbox data that we analyze in Section 5, but these deviations are minor (see discussion at beginning of Section 2). In such settings, it is reasonable to expect that the advantage of incorporating substantial additional data/information outweighs the cost of these approximations/deviations. We conclude the paper with a discussion of future directions in Section 6.
@@ 2 Methodology for learning the consumer valuation distribution $F$
We start by discussing the main assumptions needed for the subsequent methodological development. We consider a setting where we have data from several independent, non-overlapping auctions for a single (homogeneous) product. As mentioned previously, we work within the IPV framework which is quite reasonable for items/products that are used for personal consumption. Similar to George and Hui (2012), we will assume that the collection of bidders who access the auction is an i.i.d. sample from the consumer population for the corresponding product, and that the collection of private product valuations of these bidders is an i.i.d. sample from the consumer valuation distribution $F$. As stated in the introduction, it can be shown that the dominant strategy for a bidder in a second-price auction is to bid his/her true valuation. Under this strategy, any consumer who accesses the auction would bid his/her valuation with no need for multiple bidding, and we assume this behavior. We do note that in practice, some consumers do not follow this strategy. For example, eBay (1995) provides an option called proxy bidding or automatic bidding which allows the computer to automatically place multiple incremental bids below a cutoff price on behalf of the consumer (see also Ockenfels and Roth (2006); P. Bajari and A. Hortacsu (2003)). Since George and Hui (2012) only use the final selling price, they use a weaker assumption which allows for multiple bidding and stipulates that a consumer bids his/her true valuation sometime before the end of the auction.
Our final assumption is regarding the arrival mechanism of bidders in the auction. We assume that consumers arrive at/access the auction according to a Poisson process with constant rate $\lambda$. Again, a rational bidder (based on economic theory) in a second-price auction should be indifferent to the timing of his/her bid (see Milgrom (2004); Barbaro and Bracht (2021)), and this assumption makes sense in such settings. Again, we note that late-bidding (sniping) has been observed in some eBay auctions (see Bose and Daripa (2017); Barbaro and Bracht (2021)).
@@ 2.1 Joint density of the observed data in a single second price auction
Consider a given (single) second price auction with reserve price $r$. Hence, the initial selling price, denoted by $X_{0}$, is equal to $r$. The first time a consumer with bid value greater than $r$ arrives at the auction, that bid is placed but the current selling price remains $r$. Subsequently, the current selling price (standing price) changes whenever a bid greater than the existing selling price is placed. Let $M$ denote the number of times the selling price changes throughout the course of the auction. When $M>0$, let $\{X_{i}\}_{i=1}^{M}$ denote the sequence of current selling prices observed throughout the course of the auction, with $X_{i}$ denoting the new selling/standing price after the $i^{th}$ change. When $M>0$, let $T_{i}$ denote the intermediate time between the $i^{th}$ and $(i+1)^{th}$ changes in the selling/standing price for $0\leq i\leq M-1$. In particular, it follows that when $M>0$, $T_{0}$ denotes the waiting time from the start of the auction until the moment when for the the first time, the selling price changes to a higher value than the reserve price $r$. When $M=0$, we define $T_{0}=\tau$. Finally, let $O$ be a binary random variable indicating whether the item is sold before the end of the auction, i.e., $O=1$ indicates the item is sold and $O=0$ indicates that the item is not sold. Our observed data comprises of $M$, $O$, $\{X_{i}\}_{i=1}^{M}$, and $\{T_{i}\}_{i=0}^{M-1}$.
@@
Note that there are three scenarios at the end of the auction: (a) the item is sold above the reserve price ($M>0,O=1$), (b) the item is sold at the reserve price ($M=0,O=1$), and (c) the item is not sold ($M=0,O=0$). The following lemma provides a unified formula for the joint density of the observed data encompassing all these three scenarios.
@@
where $C_{1}$ does not depend on $F$, $x_{0}$ represents the reserve price ($r$), $\lambda$ denotes the constant rate of arrival of the bidders throughout the course of the auction, and $f$ represents the density function corresponding to $F$.

### v3

# A semi-parametric approach for estimating consumer valuation distributions using second price auctions

@@ Abstract
We focus on online second price auctions, where bids are made sequentially, and the winning bidder pays the maximum of the second-highest bid and a seller specified starting price. For many such auctions, the seller does not see all the bids or the total number of bidders accessing the auction, and only observes the current selling prices throughout the course of the auction. We develop a novel semi-parametric approach to estimate the underlying consumer valuation distribution based on this data. Previous semi-parametric or non-parametric approaches in the literature only use the final selling price and assume knowledge of the total number of bidders. The resulting estimate, in particular, can be used by the seller to compute the optimal profit-maximizing price for the product. Our approach is free of tuning parameters, and we demonstrate its computational and statistical efficiency in a variety of simulation settings, and also on an Xbox 7-day auction dataset on eBay.
@@ 1 Introduction
In a second price auction with starting price, the product on sale is awarded to the highest bidder if the corresponding bid is higher than a seller-specified starting price $r$. The price paid by the winner is however, the maximum of the starting price and the second highest bid. These auctions have been the industry standard for a long time, and are attractive to sellers as they induce the bidders to bid their true “private value” for the product, i.e., the maximum price they wish to pay for it. While some platforms have recently moved to first price auctions, the second price auction is still widely used on E-commerce platforms such as eBay, Rokt, and online ad exchanges such as Xandr. The analysis of data obtained from these auctions presents unique challenges. For a clear understanding of these challenges, we first discuss in detail the auction framework, the observed data and the quantity of interest that we want to estimate/extract.
@@
Auction framework: We consider an auction setting where a single product is on sale for a fixed time window $0,\tau]$. The seller sets the starting price $r$, which is used as the current selling price at time 0. Any bidder who arrives subsequently is eligible to place a bid only if his/her bid value is higher than the current selling price at that time. If the bid is placed, the current selling price is updated to the second-highest bid value among the set of all placed bids up to that time, including this latest bid (the starting price is also treated as a placed bid).  [^1] Typically, a small increment (e.g. $\mathdollar 0.01$) is also added to the second highest bid, but this insignificant increment is unlikely to influence bidder’s behaviour, and we ignore it in our analysis for ease of exposition. For example, suppose the current selling price at a given time is $\mathdollar 4$ and the highest placed bid value up to that time is $\mathdollar 5$. If a bidder comes and bids $\mathdollar 3$, the bid will not be placed. If the bidder were to bid $\mathdollar 4.5$, the bid would be placed and the current selling price would be updated to $\mathdollar 4.5$ (since now the second largest placed bid is $\mathdollar 4.5$). If the bidder were to bid $\mathdollar 6$, the bid would be placed and the current selling price would be updated to $\mathdollar 5$. At the end of the auction period, if no bid above the starting price is placed, the item goes unsold, otherwise it is sold to the highest bidder at the selling price at time $\tau$. This final selling price is the second highest placed bid (including the starting price) throughout the course of the auction.
@@
Figure 1: An illustration of a single second price auction. True bid values are generated from a Pareto distribution and starting price is at $\mathdollar 2$. Blue vertical lines are the bid values, black horizontal lines are the current selling prices, and black dots are the time points when the bids are made.
@@
An illustration of a single second price auction: Figure 1 provides a concrete illustration of how a second price auction works. The data for this auction has been simulated from a setting where the bids follow a Pareto distribution with location parameter $3$ and dispersion parameter $100$. The waiting times between bids are generated from an exponential distribution with rate parameter $1$. In total, $100$ bids are generated in a time period $\tau$ of around $115$ minutes. The starting price for the auction is $\mathdollar 2$. In Figure 1, the bid values and the current selling prices during the course of the auction are represented by the blue vertical lines and the black horizontal lines, respectively. The black dots on the black horizontal lines represent the time points (in minutes) of the $100$ bids. As can be seen from Figure 1, the initial selling price is equal to the starting price ($\mathdollar 2$). We have the first bid of around $\mathdollar 8.05$ at $0.55$ minutes. Since it’s higher than the starting price of $\mathdollar 2$, this bid will be placed, but the starting price still remains the current selling price at $0.55$ minutes. Recall that arriving consumers only see the current selling price, and hence a consumer with a $\mathdollar 5.09$ valuation would be both willing and eligible to place their bid when the current selling price is $\mathdollar 2$. Hence, the second bid of $\mathdollar 5.09$ occurring at $5.96$ minutes ($5.41$ minutes after the first bid’s occurrence) is placed, and subsequently the current selling price jumps to $\mathdollar 5.09$ as it’s the second highest value among the starting price ($\mathdollar 2$) and the two existing placed bids ($\mathdollar 8.05$ and $\mathdollar 5.09$). We don’t observe any jumps in the current selling price for the next few bids as they all are less than the current selling price of $\mathdollar 5.09$ (and hence are not eligible to be placed). Then we see another jump at $9.65$ minutes, where a bid of $\mathdollar 12.82$ which makes the current selling price jump to $\mathdollar 8.05$. The next few bids again happen to be less than the current selling price ($\mathdollar 8.05$). At around $24$ minutes, we see a last jump in the current selling price to $\mathdollar 10.14$ (based on a new bid of $\mathdollar 10.14$ which exceeds $\mathdollar 8.05$). The subsequent bids are all less than $\mathdollar 10.14$, it remains the current selling price throughout the rest of the auction period. This can be observed through the flat horizontal black line at $\mathdollar 10.14$ in the time period of $(24.7,114.43)$ minutes. The final selling price for the auction is therefore $\mathdollar 10.14$. The observed data for the above auction is the sequence of current selling prices given by $(\mathdollar 2,\mathdollar 5.09,\mathdollar 8.05,\mathdollar 10.14)$ and the sequence of times at which there was a change in the current selling price, given by $(5.96,9.65,24)$.
@@
Observed data: The observed data is the sequence of current selling price values (also sometimes referred to as the standing price) throughout the course of the auction, and the times at which there is a change in the selling/standing price. Typically, such data is available for multiple auctions of the same product. For example, in Section 5, we analyze data with current selling prices for $93$ different eBay $7$-day auctions for Xbox. A key observation to make here is that consumers who access the auction but have bids which are less than the current selling price (standing price) are not allowed to place their bids. In other words, instead of observing the bids of all the customers who access the auction, we only observe the running second largest value of such bids.
@@
Figure 2: An illustration of 4 separate second price auctions for a given product. For each auction, true bid values are generated from a discrete Uniform distribution and starting price is respectively set to $\mathdollar 10,\mathdollar 5,\mathdollar 13,\mathdollar 17$. Blue vertical lines are the bid values, black horizontal lines are the current selling prices, and black dots are the time points when the bids are made.
@@
An illustration of multiple second price auctions: Figure 2 provides a concrete illustration of a setting with multiple (four) second price auctions for the same product. The data for these auctions has been simulated from a setting where the bids follow a discrete uniform distribution with lower bound 1 and upper bound 25. The waiting times between bids are generated from an exponential distribution with rate parameter $1$, and the time period for each auction is 10 days. The starting prices for the four auctions, respectively, are $\mathdollar 10,\mathdollar 5,\mathdollar 13,\mathdollar 17$. Just like in Figure 1, the bid values and the current selling prices during the course of each auction are represented by the blue vertical lines and the black horizontal lines, respectively. The black dots on the black horizontal lines represent the time points (in days) of the bids. We can see that for auctions in the upper left and upper right, the corresponding item is sold above the starting price (at final selling price of $\mathdollar 19$ and $\mathdollar 25$ respectively). On the other hand, for the auction in the lower left, the item is sold at the starting price, since there is only one bid placed above/at the starting price. Finally, for the auction in the lower right, the item is not sold since no bid is placed above or at the starting price. We will keep coming back to this integrated example to clarify and illustrate various concepts and notions that are introduced in the paper.
@@
Quantity of interest: Each bidder in the consumer population is assumed to have an independent private valuation (IPV) of the product. The IPV assumption in particular makes sense for products that are used for personal use/consumption (such as watches, jewelry, gaming equipment etc.) and is commonly used in the modeling of internet auctions (see Song (2004); Hou and Rego (2007); George and Hui (2012) and the references therein). Economic theory suggests that the dominant strategy for a bidder in a second price auction is to bid one’s true valuation (Vickrey (1961)). The quantity that we want to estimate from the above data is the distribution of the valuations of the product under consideration for the consumer population. We refer to this as the consumer valuation distribution, and denote it by $F$. As noted in George and Hui (2012), knowledge of $F$ provides the consumer demand curve for the product, and hence can be used by the seller to identify the profit-maximizing price (see the discussion in Section 4.1 from George and Hui (2012)).
@@
This paper $2^{nd}$ Semi-parametric Yes No No No
@@
Table 1 provides a comparison of the main relevant features of key related papers and the proposed methodology in the paper. As the table demonstrates (and to the best of our knowledge), none of the existing non-parametric/semi-parametric methods for second price auctions use the entire sequence of current selling price values to estimate the consumer valuation distribution $F$. This was the primary motivation for our work, as only using the final selling price (and possibly the maximum placed bid, if available) leaves out a lot of available information. Including the current selling price information throughout the course of the auction however, involves significant conceptual and methodological challenges. As observed previously, the final selling price (second largest placed bid) is also the second largest (potential or placed) bid of all the consumers who accessed the auction. Hence, under relevant assumptions (see beginning of Section 2), the final selling price can be interpreted as the second largest order statistic of i.i.d. samples from $F$. This interpretation is central to the methodology developed in George and Hui (2012) and Song (2004). However, as the authors in George and Hui (2012) point out, the second largest current selling price throughout the course of the auction is not necessarily the third largest order statistic among all (placed or potential) bids, unless some severely restrictive and unrealistic assumptions are imposed on the order in which bidders arrive in the auction. Hence, extending the methodology in George and Hui (2012) to include the entire sequence of current selling prices is not feasible; a completely new method of attack is needed.
@@
In this paper, we fill this gap in the literature and develop novel methodology for semi-parametric estimation of the consumer valuation distribution in second-price ascending bid auctions which uses the entire sequence of current selling prices. Additionally, the total number of consumers accessing the auction is not assumed to be known (unlike George and Hui (2012)), and the highest placed bid in the auction is also not assumed to be known (unlike Backus and Lewis (2025)). Incorporating the above novel features is very challenging, and the methodological development (provided in Section 2) is quite involved. The extensive simulation results in Section 4 demonstrate the significant accuracy gains in the estimation of $F$ that can be obtained by using the entire sequence of current selling prices as opposed to just the final selling price. We note that incorporating the current selling prices during the entire auction does come with a cost. In particular, two additional assumptions (compared to those in George and Hui (2012)) that need to be made about the rate of bidder arrival and number of bids made by a single consumer. We find some deviations from these two assumptions in the eBay $7$-day Xbox data that we analyze in Section 5, but these deviations are minor (see discussion at beginning of Section 2). In such settings, it is reasonable to expect that the advantage of incorporating substantial additional data/information outweighs the cost of these approximations/deviations. We conclude the paper with a discussion of future directions in Section 6.
@@ 2 Methodology for learning the consumer valuation distribution $F$
We start by discussing the main assumptions needed for the subsequent methodological development. We consider a setting where we have data from several independent, non-overlapping auctions for a single (homogeneous) product. As mentioned previously, we work within the IPV framework which is quite reasonable for items/products that are used for personal consumption. Similar to George and Hui (2012), we will assume that the collection of bidders who access the auction is an i.i.d. sample from the consumer population for the corresponding product, and that the collection of private product valuations of these bidders is an i.i.d. sample from the consumer valuation distribution $F$. As stated in the introduction, it can be shown that the dominant strategy for a bidder in a second price auction is to bid his/her true valuation. Under this strategy, any consumer who accesses the auction would bid his/her valuation with no need for multiple bidding, and we assume this behavior. We do note that in practice, some consumers do not follow this strategy. For example, eBay (1995) provides an option called proxy bidding or automatic bidding which allows the computer to automatically place multiple incremental bids below a cutoff price on behalf of the consumer (see also Ockenfels and Roth (2006); P. Bajari and A. Hortacsu (2003)). Since George and Hui (2012) only use the final selling price, they use a weaker assumption which allows for multiple bidding and stipulates that a consumer bids his/her true valuation sometime before the end of the auction.
Our final assumption is regarding the arrival mechanism of bidders in the auction. We assume that consumers arrive at/access the auction according to a Poisson process with constant rate $\lambda$. Again, a rational bidder (based on economic theory) in a second price auction should be indifferent to the timing of his/her bid (see Milgrom (2004); Barbaro and Bracht (2021)), and this assumption makes sense in such settings. Again, we note that late-bidding (sniping) has been observed in some eBay auctions (see Bose and Daripa (2017); Barbaro and Bracht (2021)).
@@ 2.1 Joint density of the observed data in a single second price auction
Consider a given (single) second price auction with starting price $r$. Hence, the initial selling price, denoted by $X_{0}$, is equal to $r$. The first time a consumer with bid value greater than $r$ arrives at the auction, that bid is placed but the current selling price remains $r$. Subsequently, the current selling price (standing price) changes whenever a bid greater than the existing selling price is placed. Let $M$ denote the number of times the selling price changes throughout the course of the auction. When $M>0$, let $\{X_{i}\}_{i=1}^{M}$ denote the sequence of current selling prices observed throughout the course of the auction, with $X_{i}$ denoting the new selling/standing price after the $i^{th}$ change. When $M>0$, let $T_{i}$ denote the intermediate time between the $i^{th}$ and $(i+1)^{th}$ changes in the selling/standing price for $0\leq i\leq M-1$. In particular, it follows that when $M>0$, $T_{0}$ denotes the waiting time from the start of the auction until the moment when for the first time, the selling price changes to a higher value than the starting price $r$. When $M=0$, we define $T_{0}=\tau$. Finally, let $O$ be a binary random variable indicating whether the item is sold before the end of the auction, i.e., $O=1$ indicates the item is sold and $O=0$ indicates that the item is not sold. Our observed data comprises of $M$, $O$, $\{X_{i}\}_{i=1}^{M}$, and $\{T_{i}\}_{i=0}^{M-1}$.
@@
Note that there are three scenarios at the end of the auction: (a) the item is sold above the starting price ($M>0,O=1$), (b) the item is sold at the starting price ($M=0,O=1$), and (c) the item is not sold ($M=0,O=0$). The following lemma provides a unified formula for the joint density of the observed data encompassing all these three scenarios.
@@
where $C_{1}$ does not depend on $F$, $x_{0}$ represents the starting price ($r$), $\lambda$ denotes the constant rate of arrival of the bidders throughout the course of the auction, and $f$ represents the density function corresponding to $F$.